# The collision and merger products of stars do not look alike: A magnetohydrodynamics comparison

### v1

@@ Abstract
A significant fraction of stars experience close interactions, including collisions resulting from gravitational encounters and mergers within close binary systems. These processes can produce more massive stars that may give rise to relatively rare objects such as blue stragglers. Distinguishing the outcomes of collisions and mergers is challenging yet essential for interpreting observations. This study utilizes the magnetohydrodynamics code AREPO to simulate collisions and mergers of $5$ to $10\,\mathrm{M}_{\odot}$ main-sequence stars, systematically comparing the properties of the resulting products. Both collisions and mergers yield more massive, strongly magnetized, rapidly and differentially rotating stars with cores enriched in hydrogen, but notable quantitative differences emerge. Merger products exhibit core hydrogen fractions up to $10\%$ higher than those of collision products. In both scenarios, turbulent mixing amplifies magnetic field energies by $9$ to $12$ orders of magnitude. However, magnetic fields in small-impact-parameter collision products display small-scale reversals that may dissipate over time, whereas merger products and large-impact-parameter collision products develop large-scale ordered, potentially long-lived magnetic fields. Additionally, only merger products display magnetically driven, bipolar outflows with radial velocities exceeding $300$ to $400\,\mathrm{km}\,\mathrm{s}^{-1}$. These distinctions may result in divergent long-term evolutionary outcomes, which warrant further investigation in future studies.
@@
journal: ApJfacilities: Compute Canada’s Niagara cluster, MPCDF’s Orion and Viper clusters, CITA’s Sunnyvale cluster, MPA’s Ada clustersoftware: AREPO (V. Springel, 2010; R. Pakmor et al., 2016; R. Weinberger et al., 2020), MESA (B. Paxton et al., 2011)
1 Introduction
@@
Only a small fraction of stars live out their lifetimes in isolation, while the rest interact with other stars. These interactions provide pathways for stars to gain mass via dynamically induced collisions and binary mergers triggered by unstable mass transfer. Such interactions play a key role in shaping the stellar demographics in clusters and producing observable transients. Therefore, studying them is vital in understanding the evolution of star clusters and the formation of objects such as blue straggler stars, X-ray binaries, cataclysmic variables, millisecond pulsars, etc.
@@
Stellar collisions are expected to occur frequently in the cores of dense star clusters where stellar densities are high enough to allow close encounters between single stars (e.g., J. G. Hills & C. A. Day, 1976; A. Sills et al., 1997, 2005). In the densest clusters, such encounters result in runaway collisions (S. F. Portegies Zwart et al., 1999, 2004; M. Freitag et al., 2006), possibly giving rise to very massive stars. Collisions also arise from binary-single and binary-binary interactions (e.g., S. Sigurdsson & E. S. Phinney, 1993; J. M. Fregeau et al., 2004; S. Chatterjee et al., 2013), potentially dominating collisions in star clusters on account of their larger cross-sections for interactions. As a result, the rates of stellar collisions are strongly correlated with binary fractions (S. Chatterjee et al., 2013). Hydrodynamics studies of collisions in single-single (e.g., J. C. Lombardi et al., 1995, 1996; A. Sills et al., 1997, 2001; M. Freitag & W. Benz, 2005; E. Glebbeek et al., 2013; T. Ryu et al., 2024a, 2025) and binary-single interactions (J. C. Lombardi et al., 2003; E. Gaburov et al., 2010; T. Ryu et al., 2022, 2023a, 2023b, 2023a, 2024b) have shown that the resulting products are rejuvenated due to chemical mixing, puffed up, and fast rotators. Further evolution using stellar evolution codes (e.g., A. Sills et al., 1997, 2001; E. Glebbeek et al., 2013) reveals that their long-term evolution can be significantly influenced by their properties right after collision. Notably, while T. Ryu et al. (2025) conducted a magnetohydrodynamics (MHD) study of collisions of low-mass main-sequence (MS) stars, MHD simulations of massive star collisions remain largely unexplored.
@@
Close binary stars can undergo mass transfer interactions when one of them overfills their Roche lobe (e.g., S. H. Lubow & F. H. Shu, 1975; K. Sen et al., 2022; see P. Marchant & J. Bodensteiner, 2024 for recent review) and can eventually (e.g., C. A. Nelson & P. P. Eggleton, 2001; X. Chen & Z. Han, 2008) merge if mass transfer becomes unstable (e.g., G. H. Darwin, 1879). Mergers can also occur in contact binary systems (e.g., J. Henneco et al., 2024) and are observed as luminous red novae (e.g., N. Soker & R. Tylenda, 2003; R. Tylenda et al., 2011). Massive contact binaries (e.g., L. R. Penny et al., 2008; M. Kennedy et al., 2010; P. Mayer et al., 2013; O. Öztürk et al., 2014; B. Yaşarsoy & K. Yakut, 2014; J. Lorenzo et al., 2014) are, hence, promising candidates for the mergers of massive stars. Additionally, binary hardening in clusters due to binary-single interactions (e.g., D. C. Heggie, 1975; S. Banerjee & P. Ghosh, 2006) contribute to bringing stars closer to facilitate mergers. It is challenging to model stellar mergers using hydrodynamics because the thermal (or nuclear) timescales of mass transfer prior to merger are much longer than the dynamical timescales for which such studies are suited. F. R. N. Schneider et al. (2019) conducted the only MHD study of a stellar merger of two massive MS stars by employing a scheme to artificially accelerate the mass transfer phase. Analogous to collision products, merger products are extended, longer-lived, and fast-rotating.
@@
Stellar collisions and mergers are among the proposed formation channels for blue straggler stars (BSSs; see C. Wang & T. Ryu, 2024; R. D. Mathieu & O. R. Pols, 2025 for recent reviews), those located above the MS turn-off points in the color-magnitude diagrams of open and globular star clusters (e.g., A. R. Sandage, 1953; H. L. Johnson & A. R. Sandage, 1955; E. M. Burbidge & A. Sandage, 1958). More recently, stars above the MS turnoff have also been observed in young stellar clusters (e.g., C. Li et al., 2017, 2018; A. P. Milone et al., 2023; J. Bodensteiner et al., 2021), hinting at the presence of massive BSSs. The collision channel of BSS formation is believed to contribute more in very dense globular cluster cores (e.g., M. B. Davies et al., 2004), while the binary merger channel tends to dominate in less dense open clusters (e.g., R. D. Mathieu & A. M. Geller, 2009; A. M. Geller & R. D. Mathieu, 2011). Furthermore, detailed binary evolution models suggest that mergers can play an important role in the formation of (massive) BSSs in young open clusters (e.g., C. Wang et al., 2020, 2022).
[…]
@@
Stellar mergers and collisions may also pave the way in forming stars with strong magnetic fields (e.g., L. Ferrario et al., 2009; D. T. Wickramasinghe et al., 2014; F. R. N. Schneider et al., 2016; F. R. N. Schneider, 2025), as suggested by previous MHD simulations (F. R. N. Schneider et al., 2019; T. Ryu et al., 2025). Such interactions in massive stars can explain the observations of $\sim 10\%$ of O, B and A stars having strong magnetic fields $\sim 10\,\mathrm{kG}$ (e.g. J. F. Donati & J. D. Landstreet, 2009; L. Fossati et al., 2015), many of them being chemically peculiar Ap/Bp stars (e.g., H. W. Babcock, 1947; J. D. Landstreet, 1992; M. Aurière et al., 2007). These massive magnetic stars can end their lives as magnetars, which have surface magnetic fields of $10^{13}$$10^{15}\,\mathrm{G}$ (R. C. Duncan & C. Thompson, 1992; J. S. Clark et al., 2014) and are predominantly located in regions with massive OB stars such as star-forming regions and young open clusters (e.g., S. A. Olausen & V. M. Kaspi, 2014).
@@
Given the existence of astrophysical objects expected to form through both collisions and mergers, and their anticipated similarities in mass growth and potential stellar magnetism, a critical question arises: Can we clearly distinguish the resulting products of collisions and mergers? This question is particularly relevant in interpreting the origin of observed blue stragglers. Many observational studies on blue stragglers have primarily relied on host cluster properties, such as stellar density and dynamical evolution, to infer their formation mechanisms (e.g. F. R. Ferraro et al., 2012; B. Lanzoni et al., 2016) owing to the limited theoretical understanding of the similarities and differences between the outcomes of collisions and mergers. This motivates our study, which employs moving-mesh MHD to systematically compare the mixing, rotation, structures, and magnetic fields of the resulting products of collisions and mergers of massive MS stars.
@@
This paper is organized as follows. Section 2 outlines the methods we use to generate the initial conditions and perform our MHD simulations. Section 3 presents the comparative results of collisions and mergers. Section 4 discusses the implications and summarizes the conclusions.
2 Methods
2.1 Magnetohydrodynamics (MHD)
@@
We simulate collisions and mergers of stars using the 3D MHD code AREPO (V. Springel, 2010; R. Pakmor et al., 2016; R. Weinberger et al., 2020). AREPO is an arbitrary Lagrangian-Eulerian, moving-mesh code that combines elements of the Eulerian finite-volume method and the Lagrangian smooth particle hydrodynamics method. AREPO solves the Euler MHD equations on an unstructured Voronoi mesh that moves with (roughly) the local velocity of the fluid and computes the fluxes between cell interfaces using the HLLD Riemann solver (T. Miyoshi & K. Kusano, 2005; see R. Pakmor et al., 2011 for implementation). Self-gravity is included, and the Powell cleaning scheme (K. G. Powell et al., 1999; see R. Pakmor & V. Springel, 2013 for implementation) is employed to control the magnetic field divergence. Refinement is triggered when adjacent cells have volume differences of more than a factor of ten; mass refinement is not enabled. Finally, we set the minimum gravity softening factor to be one-tenth the size of the smallest cell. Several studies have employed the AREPO code to simulate various phenomena involving stars and compact objects, e.g., F. R. N. Schneider et al. (2019); P. Vynatheya et al. (2024); T. Ryu et al. (2024a, 2025).
2.2 Stellar models
@@
Figure 1: Initial density and temperature profiles of $10\,\mathrm{Myr}$-old MS stars of different masses, corresponding to Table 1. The thinner dashed and the thicker solid lines represent the 1D profiles from MESA and 3D shell-averaged and relaxed profiles from AREPO, respectively. The AREPO profiles diverge from their MESA counterparts at densities of $\lesssim 10^{-4}\,\mathrm{g}\,\mathrm{cm}^{3}\,\mathrm{s}^{-1}$ and temperatures of $\lesssim 10^{6.25}\,\mathrm{K}$.
@@
We employ the 1D stellar evolution code MESA (B. Paxton et al., 2011) to generate non-rotating stellar models of $10\,\mathrm{Myr}$-old MS stars of metallicities $Z_{\star}=0.02$ and masses ranging from $5\,\mathrm{M}_{\odot}$ to $10\,\mathrm{M}_{\odot}$. These parameter ranges probe the early evolution of stars in young stellar clusters. Following P. Vynatheya et al. (2024), we convert these 1D profiles of densities, pressures, internal energies, and chemical abundances to 3D AREPO cells using the HEALPix-shell-layering procedure (K. M. Górski et al., 2005) from S. T. Ohlmann et al. (2017), and calculate the thermodynamic variables of each cell from the Helmholtz equation of state (F. X. Timmes & F. D. Swesty, 2000). We then relax each of the AREPO stars in isolation for five stellar dynamical timescales $t_{\mathrm{dyn,\star}}=(r_{\star}^{3}/Gm_{\star})^{0.5}$, where $m_{\star}$ and $r_{\star}$ are the star’s mass and radius, respectively, to ensure that they reach a steady state. Table 1 shows the MESA-derived parameters of the stellar models we consider, with the stellar thermal timescale given by $t_{\mathrm{th,\star}}=Gm_{\star}^{2}/r_{\star}L_{\star}$, where $L_{\star}$ is the star’s luminosity provided by MESA, and the MS lifetime $t_{\mathrm{MS,\star}}$ representing the age when the core hydrogen fraction $X_{\mathrm{core}}<10^{-3}$.
@@
Intermediate and high-mass stars, which have convective cores and radiative envelopes, are observed to have a wide range of magnetic field strengths (e.g. J. F. Donati & J. D. Landstreet, 2009; R. Walder et al., 2012), with only a small minority having strong magnetic fields. Given the uncertainties in field configurations, we initialize simple dipole magnetic fields $\mathbf{B}$ centered around each star:
$\mathbf{B}(\mathbf{r})=\frac{B_{\mathrm{surf}}}{2}\frac{3\mathbf{\hat{r}}(\mathbf{\hat{r}}\cdot\mathbf{\hat{m}})-\mathbf{\hat{m}}}{(|\mathbf{r}|/r_{\star})^{3}}$ (1)
Here, $\mathbf{r}$ is the radius vector with respect to the star’s center of mass, $B_{\mathrm{surf}}$ is the surface magnetic field at the poles, and $\mathbf{\hat{m}}$ is the unit vector along the dipole axis. We choose $B_{\mathrm{surf}}=1\,\mathrm{G}$, and $\mathbf{\hat{m}}$ to be perpendicular to the plane of collision/merger. Furthermore, we limit the initial magnetic field strength to $100\,B_{\mathrm{surf}}$ to avoid very large field strengths at the core. The impact of the initial field configuration and orientation is expected to be small (see T. Ryu et al., 2025) because of the turbulent motion of the stellar material (see Section 3) during coalescence (henceforth used to collectively refer to both collisions and mergers).
We perform resolution tests to ensure that our chosen cell resolutions $N_{\mathrm{cells}}$ produce convergent results of magnetic field amplification in collisions and mergers. Checking for convergence in the magnetic energy $E_{\mathrm{B}}=\sum(|\mathbf{B}|^{2}V)/8\pi$ (Gaussian units), where $V$ is the cell volume, we opt for a final resolution of $\sim 2\times 10^{5}$ cells per $\mathrm{M}_{\odot}$ [^1] Approximate because the refinement of cells in AREPO results in a slight increase in $N_{\mathrm{cells}}$ every time step. (see the second column of Table 1). For this resolution, the numerical Reynolds number (Equation 2 in R. Pakmor et al., 2024a) $\mathrm{Re}_{\mathrm{num}}\approx 3\mathcal{L}/d_{\mathrm{cell}}\sim 700$$1000$. Here, $\mathcal{L}\sim 0.5\,r_{\mathrm{tot}}$, with $r_{\mathrm{tot}}=r_{\star 1}+r_{\star 2}$, is the turbulent injection scale and $d_{\mathrm{cell}}$ is the average diameter of the 100 smallest AREPO cells. $\mathrm{Re}_{\mathrm{num}}\gtrsim 100$ is necessary for the magnetic field amplification not to be suppressed by numerical dissipation.
@@
2.3 Initial conditions
With relaxed stellar models in hand, we run a suite of simulations of collisions and mergers. In addition to the MHD quantities, we follow the advection of two passive scalars $K_{1}$ and $K_{2}$ corresponding to the fluid cells of the two stars. $K_{1}$ and $K_{2}$ are scalar quantities initialized as ‘1’ for the cells that make up the heavier and the lighter stars, respectively, and ‘0’ for all other cells. We define an additional passive scalar $K$ as a combination of $K_{1}$ and $K_{2}$:

### v2

@@ Abstract
A significant fraction of stars experience close interactions, including collisions resulting from gravitational encounters and mergers within close binary systems. These processes can produce more massive stars that may give rise to relatively rare objects such as blue stragglers. Distinguishing the outcomes of collisions and mergers is challenging yet essential for interpreting observations. This study utilizes the magnetohydrodynamics code AREPO to simulate collisions and mergers of $5$ to $10\,\mathrm{M}_{\odot}$ main-sequence stars, systematically comparing the properties of the resulting products. Both collisions and mergers yield more massive, strongly magnetized, rapidly and differentially rotating stars with cores enriched in hydrogen, but notable quantitative differences emerge. Merger products exhibit core hydrogen fractions up to $10\%$ higher than those of collision products. In both scenarios, turbulent mixing amplifies magnetic field energies by $9$ to $12$ orders of magnitude. However, magnetic fields in small-impact-parameter collision products display small-scale reversals that may dissipate over time, whereas merger products and large-impact-parameter collision products develop large-scale ordered, potentially long-lived magnetic fields. Additionally, only merger products display magnetically driven, bipolar outflows with radial velocities up to $300\,\mathrm{km}\,\mathrm{s}^{-1}$. These distinctions may result in different long-term evolutionary outcomes, which warrant further investigation in future studies.
@@
journal: ApJfacilities: DRAC: Niagara clustersoftware: AREPO (Springel, 2010; Pakmor et al., 2016; Weinberger et al., 2020), MESA (Paxton et al., 2011)
I Introduction
@@
Only some stars live out their lifetimes in isolation, while the rest interact with other stars (e.g., Lada and Lada, 2003; Duchêne and Kraus, 2013). These interactions provide pathways for stars to gain mass via dynamically induced collisions (e.g., Hills and Day, 1976; Schneider, 2025) and binary mergers triggered by unstable mass transfer (e.g., Lubow and Shu, 1975; Nelson and Eggleton, 2001; Schneider, 2025). Such interactions play a key role in shaping the stellar demographics in clusters and producing observable transients. Therefore, studying them is vital in understanding the evolution of star clusters and the formation of objects such as blue straggler stars (BSSs; e.g., Sandage, 1953; Mathieu and Pols, 2025), X-ray binaries (e.g., Verbunt, 1993; Tauris and van den Heuvel, 2006), millisecond pulsars (e.g., Radhakrishnan and Srinivasan, 1982; Bhattacharya and van den Heuvel, 1991), thermonuclear supernova progenitors (e.g., Whelan and Iben, 1973; Webbink, 1984; Ruiter and Seitenzahl, 2025), gravitational-wave progenitors (e.g., Postnov and Yungelson, 2014; Tauris et al., 2017; Mandel and Farmer, 2022), etc.
@@
Stellar collisions are expected to occur frequently in the cores of dense star clusters where stellar densities are high enough to allow close encounters between single stars (e.g., Hills and Day, 1976; Sills et al., 1997, 2005). In the densest clusters, such encounters result in runaway collisions (Portegies Zwart et al., 1999, 2004; Freitag et al., 2006), possibly giving rise to very massive stars. Collisions also arise from binary-single and binary-binary interactions (e.g., Sigurdsson and Phinney, 1993; Fregeau et al., 2004; Chatterjee et al., 2013), potentially dominating collisions in star clusters on account of their larger cross sections for interactions. As a result, the rates of stellar collisions are strongly correlated with binary fractions (Chatterjee et al., 2013). Hydrodynamics studies of collisions in single-single (e.g., Lombardi et al., 1995, 1996; Sills et al., 1997, 2001; Freitag and Benz, 2005; Glebbeek et al., 2013; Ryu et al., 2024a, 2025) and binary-single interactions (Lombardi et al., 2003; Gaburov et al., 2010; Ryu et al., 2022, 2023a, 2023b, 2024b) have shown that the resulting products are rejuvenated owing to chemical mixing, puffed up, and fast rotators. Further evolution using stellar evolution codes (e.g., Sills et al., 1997, 2001; Glebbeek et al., 2013) reveals that their long-term evolution can be significantly influenced by their properties right after collision. Notably, while Ryu et al. (2025) conducted a magnetohydrodynamics (MHD) study of collisions of low-mass main-sequence (MS) stars, MHD simulations of massive star collisions remain largely unexplored.
@@
Close binary stars can undergo mass transfer interactions when one of them overfills their Roche lobe (e.g., Lubow and Shu, 1975; Sen et al., 2022; see Marchant and Bodensteiner, 2024 for a recent review) and can eventually merge (e.g., Nelson and Eggleton, 2001; Chen and Han, 2008) if mass transfer becomes unstable from L2-overflow (e.g., Shu et al., 1979; Pribulla, 1998; Marchant et al., 2021) or as a result of tidal instability (e.g., Darwin, 1879; Counselman, 1973). Mergers are preceded by a contact binary phase (e.g., Henneco et al., 2024) and may be observed as luminous red novae (e.g., Soker and Tylenda, 2003; Tylenda et al., 2011). Massive contact binaries (e.g., Penny et al., 2008; Kennedy et al., 2010; Mayer et al., 2013; Lorenzo et al., 2014; Öztürk et al., 2014; Yaşarsoy and Yakut, 2014) are, hence, promising candidates for the mergers of massive stars. Additionally, binary hardening in clusters due to binary-single interactions (e.g., Heggie, 1975; Banerjee and Ghosh, 2006) contributes to bringing stars closer, facilitating mergers. It is challenging to model stellar mergers using hydrodynamics because the thermal (or nuclear) timescales of mass transfer prior to merger are much longer than the dynamical timescales for which such studies are suited; nevertheless, hydrodynamics simulations of mergers of stars have been performed (e.g., Rasio and Shapiro, 1995; Nandez et al., 2014; Motl et al., 2017; Shiber et al., 2024). Schneider et al. (2019) conducted the only MHD study of a stellar merger of two massive MS stars by employing a scheme to artificially accelerate the mass transfer phase. Analogous to collision products, merger products are extended, longer-lived, and fast-rotating.
@@
Stellar collisions and mergers are among the proposed formation channels for BSSs (see Wang and Ryu, 2024; Mathieu and Pols, 2025 for recent reviews), those located above the MS turn-off points in the color-magnitude diagrams of open and globular star clusters (e.g., Sandage, 1953; Johnson and Sandage, 1955; Burbidge and Sandage, 1958). More recently, stars above the MS turnoff have also been observed in young stellar clusters (e.g., Li et al., 2017, 2018; Bodensteiner et al., 2021; Milone et al., 2023), hinting at the presence of massive BSSs. The collision channel of BSS formation is believed to contribute more in very dense globular cluster cores (e.g., Davies et al., 2004), while the binary merger channel tends to dominate in less dense open clusters (e.g., Mathieu and Geller, 2009; Geller and Mathieu, 2011). Furthermore, detailed binary evolution models suggest that mergers can play an important role in the formation of (massive) BSSs in young open clusters (e.g., Wang et al., 2020, 2022).
In the field, stellar collisions and mergers are expected to occur frequently in triple- and quadruple-star systems (e.g., Hamers et al., 2022; Preece et al., 2024; Shariat et al., 2025; Kummer et al., 2025), where more than $50\%$ of all stars, especially massive ones, reside (e.g., Duchêne and Kraus, 2013; Moe and Di Stefano, 2017; Offner et al., 2023). For example, Preece et al. (2024) found that $30\%$$40\%$ of massive multiple-star systems undergo these interactions, primarily facilitated by significant orbital perturbations arising from (von Zeipel)-Lidov-Kozai oscillations (von Zeipel, 1910; Lidov, 1962; Kozai, 1962; see Naoz, 2016 for review). However, these field stars are challenging to identify observationally, since, unlike in star clusters, there are no other stars around to compare relative ages.
@@
Stellar mergers and collisions may also pave the way to forming stars with strong magnetic fields (e.g., Ferrario et al., 2009; Wickramasinghe et al., 2014; Schneider et al., 2016; Schneider, 2025), as suggested by previous MHD simulations (Schneider et al., 2019; Ryu et al., 2025). Such interactions in massive stars can explain the observations of $\sim 10\%$ of O, B, and A stars having strong magnetic fields $\sim 10\,\mathrm{kG}$ (e.g., Donati and Landstreet, 2009; Fossati et al., 2015), many of them being chemically peculiar Ap/Bp stars (e.g., Babcock, 1947; Landstreet, 1992; Aurière et al., 2007). These massive magnetic stars can end their lives as magnetars, which have surface magnetic fields of $10^{13}$$10^{15}\,\mathrm{G}$ (Duncan and Thompson, 1992; Clark et al., 2014) and are predominantly located in regions with massive OB stars such as star-forming regions and young open clusters (e.g., Olausen and Kaspi, 2014).
@@
Given the existence of astrophysical objects expected to form through both collisions and mergers and their anticipated similarities in mass growth and potential stellar magnetism, a critical question arises: can we clearly distinguish the resulting products of collisions and mergers? This question is particularly relevant in interpreting the origin of observed BSSs. Many observational studies on BSSs have primarily relied on host cluster properties, such as stellar density and dynamical evolution, to infer their formation mechanisms (e.g., Ferraro et al., 2012; Lanzoni et al., 2016) owing to the limited theoretical understanding of the similarities and differences between the outcomes of collisions and mergers. This motivates our study, which employs moving-mesh MHD to systematically compare the mixing, rotation, structures, and magnetic fields of the resulting products of collisions and mergers of massive MS stars.
@@
This paper is organized as follows. Section II outlines the methods we use to generate the initial conditions and perform our MHD simulations. Section III presents the comparative results of collisions and mergers. Section IV discusses the implications and summarizes the conclusions.
II Methods
II.1 Magnetohydrodynamics
We simulate collisions and mergers of stars using the 3D MHD code AREPO (Springel, 2010; Pakmor et al., 2016; Weinberger et al., 2020). AREPO is an arbitrary Lagrangian-Eulerian, moving-mesh code that combines elements of the Eulerian finite-volume method and the Lagrangian smooth particle hydrodynamics method. Several studies have employed the AREPO code to simulate phenomena involving stars, such as tidal disruption events (e.g., Ryu et al., 2023a, b, 2024b; Vynatheya et al., 2024), collisions of stars (e.g., Ryu et al., 2024a, 2025), mergers of stars (e.g., Schneider et al., 2019) and compact objects (e.g., Pakmor et al., 2013, 2021; Gronow et al., 2021; Burmester et al., 2023; Lioutas et al., 2024; Morán-Fraile et al., 2024; Glanz et al., 2025), and common envelope evolution (e.g., Ohlmann et al., 2016b, a; Sand et al., 2020; Kramer et al., 2020; Ondratschek et al., 2022; Vetter et al., 2025).
@@ II.1 Magnetohydrodynamics
AREPO solves the Euler MHD equations on an unstructured Voronoi mesh that moves with (roughly) the local velocity of the fluid and computes the fluxes between cell interfaces using the HLLD Riemann solver (Miyoshi and Kusano, 2005; see Pakmor et al., 2011 for implementation). Self-gravity is included, and the Powell cleaning scheme (Powell et al., 1999; see Pakmor and Springel, 2013 for implementation) is employed to control the magnetic field divergence. Refinement is triggered when adjacent cells have volume differences of more than a factor of ten; mass refinement is not enabled. Finally, we set the minimum gravity softening factor to be one-tenth the size of the smallest cell.
II.2 Stellar models
@@
Figure 1: Initial density and temperature profiles of $10\,\mathrm{Myr}$-old MS stars of different masses, corresponding to Table 1. The thinner dashed and thicker solid lines represent the 1D profiles from MESA and 3D shell-averaged and relaxed profiles from AREPO, respectively. The AREPO profiles diverge from their MESA counterparts at densities of $\lesssim 10^{-4}\,\mathrm{g}\,\mathrm{cm}^{3}\,\mathrm{s}^{-1}$ and temperatures of $\lesssim 10^{6.25}\,\mathrm{K}$.
@@
We employ the 1D stellar evolution code MESA (Paxton et al., 2011) to generate nonrotating stellar models of $10\,\mathrm{Myr}$-old MS stars of metallicities $Z_{\star}=0.02$ and masses ranging from $5\,\mathrm{M}_{\odot}$ to $0\,\mathrm{M}_{\odot}11$1The MESA star inlists are available at 10.5281/zenodo.18249784 (catalog doi: 10.5281/zenodo.18249784).. These parameter ranges probe the early evolution of stars in young stellar clusters. Following Vynatheya et al. (2024), we convert these 1D profiles of densities, pressures, internal energies, and chemical abundances to 3D AREPO cells using the HEALPix-shell-layering procedure (Górski et al., 2005) from Ohlmann et al. (2017) and calculate the thermodynamic variables of each cell from the Helmholtz equation of state (Timmes and Swesty, 2000). We then relax each of the AREPO stars in isolation for five stellar dynamical timescales $t_{\mathrm{dyn,\star}}=(r_{\star}^{3}/Gm_{\star})^{0.5}$, where $m_{\star}$ and $r_{\star}$ are the star’s mass and radius, respectively, to ensure that they reach a steady state. Table 1 shows the MESA-derived parameters of the stellar models we consider, with the stellar thermal timescale given by $t_{\mathrm{th,\star}}=Gm_{\star}^{2}/r_{\star}L_{\star}$, where $L_{\star}$ is the star’s luminosity provided by MESA, and the MS lifetime $t_{\mathrm{MS,\star}}$ representing the age when the core hydrogen fraction $X_{\mathrm{core}}<10^{-3}$.
@@
Intermediate- and high-mass stars, which have convective cores and radiative envelopes, are observed to have a wide range of magnetic field strengths (e.g., Donati and Landstreet, 2009; Walder et al., 2012), with only a small minority having strong magnetic fields. Given the uncertainties in field configurations, we initialize simple dipole magnetic fields $\bm{B}$ centered around each star:
$\bm{B}(\bm{r})=\frac{B_{\mathrm{surf}}}{2}\frac{3\bm{\hat{r}}(\bm{\hat{r}}\cdot\bm{\hat{m}})-\bm{\hat{m}}}{(|\bm{r}|/r_{\star})^{3}}$ (1)
Here, $\bm{r}$ is the radius vector with respect to the star’s center of mass (CoM), $B_{\mathrm{surf}}$ is the surface magnetic field at the poles, and $\bm{\hat{m}}$ is the unit vector along the dipole axis. We choose $B_{\mathrm{surf}}=1\,\mathrm{G}$, and $\bm{\hat{m}}$ to be perpendicular to the orbital plane. Furthermore, we limit the initial magnetic field strength to $100\,B_{\mathrm{surf}}$ to avoid very large field strengths at the core. The impact of the initial field configuration and orientation is expected to be small (see Ryu et al., 2025) because of the turbulent motion of the stellar material (see Section III) during coalescence (henceforth used to collectively refer to both collisions and mergers).
We perform resolution tests to ensure that our chosen cell resolutions $N_{\mathrm{cells}}$ produce convergent results of magnetic field amplification in collisions and mergers. Checking for convergence in the magnetic energy $E_{\mathrm{B}}=\sum_{\mathrm{cells}}(|\bm{B}_{\mathrm{cell}}|^{2}V_{\mathrm{cell}})/8\pi$ (Gaussian units), where $V_{\mathrm{cell}}$ is the cell volume, we opt for a final resolution of $\sim 2\times 10^{5}$ cells per $\mathrm{M}_{\odot}$ [^2] Approximate because the refinement of cells in AREPO results in a slight increase in $N_{\mathrm{cells}}$ every time step. (see the second column of Table 1). For this resolution, the numerical Reynolds number (Equation 2 in Pakmor et al., 2024a) $\mathrm{Re}_{\mathrm{num}}\approx 3\mathcal{L}/d_{\mathrm{cell}}\sim 700$$1000$. Here, $\mathcal{L}\sim 0.5\,r_{\mathrm{tot}}$, with $r_{\mathrm{tot}}=r_{\star 1}+r_{\star 2}$, is the turbulent injection scale and $d_{\mathrm{cell}}$ is the average diameter of the 100 smallest AREPO cells. $\mathrm{Re}_{\mathrm{num}}\gtrsim 100$ is necessary for the magnetic field amplification not to be suppressed by numerical dissipation.
@@
II.3 Initial conditions
With relaxed stellar models in hand, we run a suite of simulations of collisions and mergers. In addition to the MHD quantities, we follow the advection of two passive scalars $K_{1}$ and $K_{2}$ corresponding to the fluid cells of the two stars. $K_{1}$ and $K_{2}$ are scalar quantities initialized as ‘1’ for the cells that make up the heavier and lighter stars, respectively, and ‘0’ for all other cells. We define an additional passive scalar $K$ as a combination of $K_{1}$ and $K_{2}$: